-- Public.MarkWard - 05 Jul 2017


  • Built script to extract anode and ground baselines based on the first few thousand samples (between the start -400ns, should probably be tightened up)
  • Fits to baselines give parameters to perform data quality cuts, tell us which events in a run can be discarded.
  • Baseline_Vs_Trigger.png
  • Baseline_Sigma.png
  • TriggerRates.png
  • Lab preparation, Zach and I mocked up an abomination of roughly the right maximum dimentions of the chamber and successfully passed it through the corridors to the lab.
  • Turns out the roof of the clean tent is very heavy, thanks to everyone involved, minor emergency averted but the method of supporting the roof has to be much much stronger
  • Probably use a shoring prop from a builders supplier if we cant come up with a better option
  • Upstairs lab cleanup - We tidied up DARCO, The stand now has the following
    • 1 power plug that feeds
    • Camera power plug is independant
    • High pressure gauge is independant
  • It can be wheeled anywhere around the lab a requires 3 sources of power rather than how it was before, using 3 power bars with a squid work of cables. Much tidier and organised now.


  • Lab Clean and preparation for recieving HPTPC
  • Added supporting option for removal of the clean room bottom right foot, it is a labjack attached to the bottom left leg of the annex via 80-20, Iv not removed the leg but it could be removed at any time to test the loading
  • HPTPC delivery delayed -> Joe has the full details
  • Fixed (fingers crossed) the roof of the clean tent annex. Adhesive tape leached the plasticiser from the PVC and turned into goop, it isn't perfect but it is better than it being wide open!
  • Clean Room (based on it having a roof for one day) has reduced the dust levels by roughly factor of 2 over the outside air. Could probably be better once all the roof is fixed.
  • Spent time working with the DMTPC software, trying to familiarise myself with the functions and what they do.


  • Added light tightening measure around camera case
    • Black insulation tape, cut around the USB and Power only
    • Aluminium tape, same treatment
  • See Will's PRD meeting talk for the results -> LightLeakImages.pdf
  • Changed cathode feedthrough
  • 2 steps forward, one step back - the light features in the new images look like a light source on the top left, casting light on the camber. You can convince yourself you are looking at an out of focus right hand rail
  • Examined spare feedthrough - it is clearly transmitting diffuse light
  • Made a foam and tape cover for the top, light tightened the high voltage cable. This fixes the issue!
  • Voltages
    • Maximum Cathode voltage reached 2800-2900V before producing breakdown within ~few seconds or not making the ramp to the voltage
    • Stable for runs 2500-2600V
  • Have observe the cathode feedthrough producing a glow,
  • Spark rates appear higher with higher cathode voltage (no numbers, but just based observing behavour... seems like it makes sense given the charge collection should be better/more dense)
  • CathodeT_Glow.png
  • Tracks_again.png
  • Taken a batch of 40 runs - about 3000 seconds of data


  • Low amounts of updates - Sorry
  • Previous Amplitude plot appears to misrepresent the data - plotting a graph rather than histogram reveals lots of negative Y axis points aaaand oddly the population at 600goes away - need to investigate why!?!
  • Previous tasks, emphasis was given to designing new mounts for the HPTPC vessel, Can't really build a TPC without a way of mounting it.
  • Progress : Runners and rails - Done - Need review and sending to the workshop.
  • Internal_Rail_Drawing_v1.pdf
  • Field cage mounting system - Done - needs preparing into a drawing package, some screen shots
  • RailFullTop.pngRailOpenTop.png
  • RailOpenBottom.png
  • Each section bolts to the last in a string, via the 'bottom outer' 'top inner' bolting positions which allows the mount to have some rigidity
  • A captured threaded rod busing arraingment means we don't require right handed and left handed threaded sections for adjustment
  • We have 1 - 1.5 cm adjustment, we could try and design around more, requires the Field cage to be the correct diamater wink
  • Slotted holes almost identical to the current configuration for the resistive chain.
  • All mounts are identical, no top and side rail configurations, there is not enough space or need for a differential support system
  • Will prepare a drawing package
  • To do : Cathode mount, and Anode mount - Should be simpler! (should)


  • Moire pattern is visible in the pre-RHUL move and install
  • DSC05271.JPG
  • The patterning also appears similar in structure to that of the ion gauge picture, Top Left looks smooth, Top Right and Bottom left look distinct, and bottom right looks like there is a fine grid
  • Looking back at our 40mbar CF4 data, I found that we had very few triggers, example, 700V data, with 40 1s exposures had a total of 10 triggers, that where actually spread over 3 events rather than a steady rate
  • Turned out that looking at the config files, the channels where plugged into the digitizer as shown in photographs, and it appears to have been exactly reversed. Probable that the digitizer had already been disconnected and then quickly reconnected for a photo.
  • Now getting a more reasonable trigger rate given there is an Am241 source right next to one quadrant.
  • Getting practice in extracting waveforms from the data files so i can understand better how the files are structured and what data i can pull. I have taken 40 run sequences each with 20 exposures at 2 seconds per exposure. Should have a reasonable data sample to get started on the gain (if the low cathode voltage doesn't prevent that)
  • So, Good/interesting data is for 1600V Cathode - 710V Anode - 40.18mbar
  • Some examples
  • First Run of our sample - Event 1 has 16 triggers, all in the bottom left quadrant
  • Triggers_Examples_R1227005.png
  • Triggers_Example_R1227005.png
  • Rise time vs Amplitude plot - All channels? (need to check what the code is doing)
  • RisevsAmp.png
  • Cathode sparking, is definitely detectorside and not a problem with the powersupply. I switched for a different supply and it was showing evidence of sparking. It being possible to hear the breakdown around the feedthrough. We should plan on either bakeout/modification/swapping for a bigger one when we next open up.


  • Took data with waveforms, though appears not to be many events in those files,
  • Opened Chamber and Replaced Cathode cable to Kapton Insulated copper
  • Cathode HV ramp in air successful to 5KV
  • Installation of a quartz window on the left middle 4.5inch CF half nipple
  • Current Vacuum achieved is 2.7x10^-6 - an improvement, maybe previous cathode cable was outgassing significantly.
  • Installed LED light source for testing.
  • Did focus testing with top right CCD in order to get a good focus. Setting that (by eye) gives good focus is just less than 0.6m on the focus ring (DOF indicator over the . of 0.6)
  • IMG_20170807_172149.jpg
  • Matched all 4 cameras to same focus - Image below is a very lazy splicing together of the results, which are a vast improvement (originally so out of focus, you couldnt really see the anode frame)
  • All.png
  • Followed up with tests of the light source - Example with Red LED - operating at 1Hz, 20% duty, 0.7vPP, 0.35 offset
  • LightSource.png
  • LightSourceTR.png

  • Light tightness tests
  • Very preliminary though with our run config appears to be working, bias frames are taken shutter closed, proven during the focus testing
  • Setting of the shutter closed during exposures also works, proven we can run with the shutter never moving.
  • The features dicussed previously appear to be present in the bias frames (shutter closed)
  • Magnitude of the features appears to change with time, though evidence that they tend to reduce during runs (so example, 60s exposures, once 3 or so frames are taken, the general magnitude appears to reduce)
  • Still need to investigate more quantitively, though could it be a thermal effect? we want to try different run configurations, lots of shorter exposures vs long exposures, running in sequence so the bias frames are taken in similar thermal conditions etc... still under investigation.


  • Extensive voltage ramping of Anode/Cathode to gauge where we can get to reasonably
  • BL_V.gif
  • BL_I.gif
  • BL_rate.gif
Voltage 714 735 760 748
Rate at stop 0.04* 0.04 0.025 0.025*

  • All tests where stopped when rates started to rise rapidly, TR could probably be trained and pushed further
  • TL shows breakdown/sparking at much lower voltages now, possible the initial test dislodged something, might need some TLC, could also be a rise of the previous issues with that quadrant
  • Pumped down vessel over weekend, acheived better vacuum (3.3x10^-6mbar compared to 4.8x10^-6 mbar) in those two days, looks positive, and filled back up to 40mbar
  • Rammped Up all quadrants at the same time, Cathode can only acheive roughly 1700V if ramped very slowly, possible breakdown causing the trip.
  • Highest operating voltage accross the whole TPC acheived thus far is 650(TL) and 724 for all others, TR has highest spark rate, again this could be related to the top support.
  • Attached all amplifier power (longer task that you might think) and performed several ramps.
  • Moved power distribution in the lab around, allowing us to hard reset the Cathode power supply and also power on and off the amplifiers
  • Taken gain data at 670V through 720V in 10V steps, only short runs of about 300 seconds, Cathode voltage at 1500V
  • Still need to look at the data, no results.

  • Door widening is complete
  • approx 4 rounds of cleaning in the lab
  • Particle counts are returning to normal in the room, inside the clean tent levels are approximately normal
  • Moved CF4 into the lab
  • Pressurised DMTPC to 30.83torr (41.1mbar)
  • Pressure.gif
  • TL Ramping
  • TL_v.gif
  • TL_i.gif
  • Final Voltage before ramp down 714.4 V, Spark rate ~ 1Hz
  • Back down to 650V - spark rate reduced to ~0.4Hz
  • Voltage Ramping of TR quadrent
  • TR_v.gif
  • TR_i.gif
  • Final voltage before ramping down 735.2
  • Spark Rate.....?


  • Prepared for installation of DMTPC field cage
    • Equipment prep and cleaning
    • Lab cleaning
    • Electrical testing - spark test of Anode
    • TR TL BR BL
      2.5kV 2.3kV 2.54kV 2.5kV
  • Installed feedthroughs for Cathode Ground and Anode
    • Opened chamber
    • Connected anode and ground through feedthroughs, debugged a fault involving a grounding amiplifier which appears not to give the correct potential between ground and field cage.
    • Solved a problem involving measuring voltage drops across 100s MOhm circuits. Needed a high impedence multimeter, was able to prove the voltage out is in line with voltage in.
    • Performed voltage ramp of the vessel
    • TR TL BR BL
      2.55kV 2.6kV 2.55kV 2.6kV
      shows improvement on all voltages achieved
    • positioned anode 48.5mm from the flange
    • Installed Field cage and Cathode onto rails, made grounding connection and high voltage connection
    • Measured 3.0 MOhm across each step and likewise voltage delivery and equal step downs between different field cage ring.
  • Vessel Closing
    • Closure of vessel was quick thanks to correct positioning of it (in height) during the first opening and closing operation.
    • Rapid evacuation to 1E-4mbar within 20mins of closure
    • Estimate 4:30hrs of vessel open exposure.
    • Current vacuum level is 8.2e-6mbar
    • value_in_chamber_pressure_bpg_tmp1500456290.gif
  • Lab Door widening
    • Workmen have put a tape tent around the entrance though it isn't perfect.
    • Room will be quite dusty!
    • We want to do a fill with some gas - investigation -> very hard to get a bottle in without basically removing a large section of sheeting at the side and introducing even more dust. Entrance zipper stops about 4-6 inches from the floor and is tricky to get a bottle over.
    • Need to investigate further today - Yesterday the presense of supports to hold the door lintel didn't make me want to swing a bottle around in that area given it was being held in place with... wet cerment/glue (ie supported completely by the floor pillars)
TPC Install Photos


  • DM-HP-TPC_Task_List-Google_Sheets.pdf: DM-HP-TPC_Task_List-Google_Sheets.pdf
  • DMTPC - Rails installed, though not fully outgased as we lost the water cooling before we could pump on things for a long time
  • Cooling is back now (2-3 day downtime)
  • TPC transfer frame completed to correct dimensions, Cathode and fieldcage installed onto the transfer frame
  • Door resizing will be started on monday, - Target - TPC will be installed in the chamber before then! GO GO GO
  • Cage.jpg:
  • Field Cage resistor chain is installed, link to cathode and link to ground to be done (but are single parts)
  • Anode has been unpacked and has been hung on the frame with supports at the correct size
  • IMG_20170711_170058.jpg:
  • Need to make modification to bolting hardware, on the top mount (second part not with the equipment is adiquetly lengthed nylon or delrin bolts)
  • Next steps will be cabling prep - on frame in air voltage tests


  • leak check on the vessel is excellent, 2.5x10^-6 mbar l / s
  • Reached 10^-9 mbar range in pressure
  • After a failed fitting of the TPC to the transfer frame, we decided to open the chamber and take field measurements and fit the rails.
  • Huge lab cleanup last week - no cardboard - everything in plastic tubs/bags
  • Monday was spent planning and cleaning components/the lab
  • Tuesday
    • Opened the chamber
    • Measured the Rail locations with respect to each other, Does not match the CAD!
    • Rail is wider at the back (slightly)
    • Bottom rails are not centred with respect to the top rail
    • TPC support frame is machiened incorrectly for the Cathode which puts it off centre.
  • Vessel was successfully closed with the rails installed
  • Pumpdown was successful - though slow - remember, plastics = water = bad, currently high 10^-6 mbar
  • Next step is to change the frame to reflect the new measurements, this will allow almost complete construction of the TPC ready for popping in the vessel.
  • Cable lengths are being determined, ordering more cable asap, just recieved specs
  • Totally just saw a supernova...
Topic attachments
I Attachment History Action Size Date Who Comment
PNGpng All.png r1 manage 2079.8 K 09 Aug 2017 - 10:35 MarkWard  
GIFgif BL_I.gif r1 manage 26.5 K 02 Aug 2017 - 10:44 MarkWard  
GIFgif BL_V.gif r1 manage 24.0 K 02 Aug 2017 - 10:41 MarkWard  
GIFgif BL_rate.gif r1 manage 18.5 K 02 Aug 2017 - 10:41 MarkWard  
PNGpng Baseline_Sigma.png r1 manage 183.4 K 20 Sep 2017 - 10:43 MarkWard  
PNGpng Baseline_Vs_Trigger.png r1 manage 176.7 K 20 Sep 2017 - 10:43 MarkWard  
JPEGjpg Cage.jpg r1 manage 2040.2 K 12 Jul 2017 - 11:21 MarkWard  
PNGpng CathodeT_Glow.png r1 manage 282.2 K 06 Sep 2017 - 10:49 MarkWard  
PDFpdf DM-HP-TPC_Task_List-Google_Sheets.pdf r1 manage 514.6 K 12 Jul 2017 - 11:08 MarkWard  
JPEGjpg DSC05271.JPG r1 manage 3470.4 K 16 Aug 2017 - 10:01 MarkWard  
JPEGjpg IMG_20170711_170058.jpg r1 manage 2140.9 K 12 Jul 2017 - 11:10 MarkWard  
JPEGjpg IMG_20170807_172149.jpg r1 manage 2334.1 K 09 Aug 2017 - 10:30 MarkWard  
PDFpdf Internal_Rail_Drawing_v1.pdf r1 manage 263.2 K 30 Aug 2017 - 10:18 MarkWard  
PDFpdf LightLeakImages.pdf r1 manage 6329.6 K 06 Sep 2017 - 10:32 MarkWard  
PNGpng LightSource.png r1 manage 72.9 K 09 Aug 2017 - 10:47 MarkWard  
PNGpng LightSourceTR.png r1 manage 79.2 K 09 Aug 2017 - 10:51 MarkWard  
GIFgif Pressure.gif r1 manage 10.5 K 26 Jul 2017 - 10:30 MarkWard  
PNGpng RailFullTop.png r1 manage 183.7 K 30 Aug 2017 - 10:25 MarkWard  
PNGpng RailOpenBottom.png r1 manage 118.0 K 30 Aug 2017 - 10:30 MarkWard  
PNGpng RailOpenTop.png r1 manage 222.2 K 30 Aug 2017 - 10:27 MarkWard  
PNGpng RisevsAmp.png r1 manage 28.0 K 16 Aug 2017 - 10:43 MarkWard  
GIFgif TL_i.gif r1 manage 10.5 K 26 Jul 2017 - 10:30 MarkWard  
GIFgif TL_v.gif r1 manage 11.8 K 26 Jul 2017 - 10:30 MarkWard  
GIFgif TR_i.gif r1 manage 13.6 K 26 Jul 2017 - 10:30 MarkWard  
GIFgif TR_v.gif r1 manage 11.3 K 26 Jul 2017 - 10:30 MarkWard  
PNGpng Tracks_again.png r1 manage 525.7 K 06 Sep 2017 - 11:00 MarkWard  
PNGpng TriggerRates.png r1 manage 84.6 K 20 Sep 2017 - 10:44 MarkWard  
PNGpng Triggers_Example_R1227005.png r1 manage 544.1 K 16 Aug 2017 - 10:35 MarkWard  
PNGpng Triggers_Examples_R1227005.png r1 manage 50.2 K 16 Aug 2017 - 10:35 MarkWard  
GIFgif value_in_chamber_pressure_bpg_tmp1500456290.gif r1 manage 9.2 K 19 Jul 2017 - 10:33 MarkWard  
Edit | Attach | Watch | Print version | History: r13 < r12 < r11 < r10 < r9 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r13 - 20 Sep 2017 - MarkWard

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback